OpenStudy (anonymous):

tan^4(x)-sec^4(x)= 1 - 2ssec^2(x)

OpenStudy (anonymous):

\[= (-1)(2\sec^2x - 1) = 1 - 2\sec^2x\]\[= (\tan^2x - \sec^2x)(\tan^2x + \sec^2x) = (\sec^2x - 1 - \sec^2x)(\sec^2x - 1 + \sec^2x)\]

OpenStudy (anonymous):

sorry for the mess + lag : = (tan^2x - sec^2x)(tan^2x + sec^2x) = ( sec^2x - 1 - sec^2x)(sec^2x -1 + sec^2x) = (-1)(2sec^2x -1) = 1 - 2sec^2x hence: sec^2x - tan^2x = 1

OpenStudy (anonymous):

\[x=n \pi \pm \pi/3\]